OpenStudy (anonymous):

Can someone help me on these two absolute value equations? 1. lr-8l=5 2. 4l2y-3l-1=11 Please show me how to solve this step-by-step.

OpenStudy (pixiedust1):

Okay, so for this question, we are trying to find the value of r. We have to flip the equation around to get the answer. 8+5=13 Therefore, r=13 Let's check our work. I13-8I=5 The I I just means that the difference between the two numbers will be positive, but since it already is, we can just ignore that. #1 is r=13 Does that make sense?

OpenStudy (anonymous):

Yes it does, thx

OpenStudy (anonymous):

Oh ya, and I am supposed to write no solution to one of the questions so I guess #2 is the one

OpenStudy (pixiedust1):

Great! Now, let's figure out #2. I'm gonna give you a link for future math problems, and you can try using it for this one too. https://www.mathway.com/

OpenStudy (anonymous):

Thanks!

OpenStudy (pixiedust1):

No solution? I'm not sure what that means :( But those calculators are SO SO SO helpful!! there are so many online. The one I gave you is half-free, but with a simple Google search, you can find many free ones.

OpenStudy (anonymous):

Oh ok, because it says solve the equation and if you don't find a solution then write no solution.

OpenStudy (pixiedust1):

Oh I get it! By the way, the answer to the second one is 3. I'll show you how to tell. I'm gonna plug in 3 for the value of y. 2*3-3=3 4*3=12 12-1=11 Therefore, y=11

OpenStudy (pixiedust1):

I hope you find that helpful! Your such a good learner that I gave you a badge!!!

OpenStudy (anonymous):

Thx!

OpenStudy (pixiedust1):

You are very welcome! Do you have any other questions today?

OpenStudy (anonymous):

Do you know solve to solve unions?

OpenStudy (anonymous):

How to solve*

OpenStudy (pixiedust1):

Hmmm... I'm not sure what a "union" is, but if I see the problem, I may recognize it (I always forget my math term rofl)

OpenStudy (anonymous):

A U C Let A={1,3,4} C={2,5,7,10} I guess the U means union

OpenStudy (pixiedust1):

Woah that's tough. I believe I know how to solve it, but I'm not certain. Therefore, I do not want to lead you down the wrong path. :(

OpenStudy (anonymous):

The second one is C upside down U D and C={2,5,7,10} and D={x l x is an odd whole number less than ten

OpenStudy (anonymous):

Ok it's ok

OpenStudy (pixiedust1):

@phi @kelliegirl33 They can help you :)

OpenStudy (anonymous):

Ok thanks

OpenStudy (pixiedust1):

Have a blessed day! :)

OpenStudy (phi):

*** A U C Let A={1,3,4} C={2,5,7,10} **** union means the set that contains the elements from both sets, but with *no duplicates* example {1,2} U { 2,3} = {1,2,3} (not {1,2,2,3} ) intersection of two sets contains only the elements found in both sets. example {1,2} \(\cap\) { 2,3} ={2}